OpenStudy (anonymous):

MC question! A person who weights 800N steps onto a scale that is on the floor of an elevator car. If the elevator accelerates upward at a rate of 4.9m/s^2, what will the scale read? A. 400N B. 800N C. 1000N D. 1200N E. 1600N - Answer's D, but how?

OpenStudy (anonymous):



OpenStudy (loser66):

upward F = m(g+a) downward F = m(g-a)

OpenStudy (loser66):

when the elevator is at rest, F = mg --> m = F/g = 800/9.8 = 81.63kg so, upward F = 81.63 (9.8 +4.9)=1199.961 N = 1200N , it's D

OpenStudy (anonymous):

Got it @Loser66! Thanks :)